OpenStudy (anonymous):

A bag contains 4 red and 12 blue marbles. What is the probability of picking out a red marble, keeping it out of the bag, and then picking out another red marble? 27. Find the 20th term of the arithmetic sequence 9, 13, 17, 21, ... 28. Find the 17th term of the arithmetic sequence 12, 3, –6, –15, ...

OpenStudy (anonymous):

@aaronq

OpenStudy (aum):

There are a total of 16 marbles and 4 of them are red. What is the probability of picking a red marble?

OpenStudy (anonymous):

4/16? @aum

OpenStudy (aum):

correct. Now if a red marble is picked and not put back we will have 3 red marbles left and a total of 15 marbles. What is the probability of picking a red marble now?

OpenStudy (anonymous):

3/15? @aum

OpenStudy (aum):

correct. The combined probability is the product of the two probabilities: 4/16 * 3/15 = 1/4 * 1/5 = 1/20

OpenStudy (anonymous):

1/20 is the answer right ? thank you so much @aum

OpenStudy (aum):

You are welcome.

OpenStudy (anonymous):

do you think you can help 27 and 28 ?? @aum